OpenStudy (anonymous):

Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this 4x – 5y = -31 -12x + 15y = 93

OpenStudy (anonymous):

are you still there??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lets start by eliminating y :\[~3~(4x ~–~ 5y ~= -31)\]\[-12x + 15y = 93\] \[12x-15y=-93\]\[-12x+15y=93\] \[0=0\] as you can see this results in 0=0 and that means that there are an 'infinite number of solutions'. \[\large:)\]

OpenStudy (anonymous):

@amber517 i hope this helps! :)

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

you're welcome...

OpenStudy (anonymous):

:D